OpenStudy (anonymous):

A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture? 15 10 12

OpenStudy (anonymous):

You have 30 pounds of chocolates - "x" pounds of the first type, and "30-x" of the second type. The combined chocolates are selling for $1.10 per pound, so the entire batch is $33 (30*1.10 == 33). If your "x" pounds is the $1.50 chocolates, you get: 1.50x + 0.90(30 - x) = 33 Then, you solve for x. Since I hate dealing with decimals, I'm going to multiply both sides by 10 first: 15x + 9(30 - x) = 330 Expanding the second term: 15x + 270 - 9x = 330 Subtracting 270 from both sides, and combining the x terms: 6x = 60 Dividing through by 6: x = 10 So, you have 10 pounds of the expensive chocolates, and 30-10=20 pounds of the cheap ones.

OpenStudy (anonymous):

so 10

OpenStudy (anonymous):

yes